### v1

@@ 2.2 Wasserstein Ambiguity Sets
Wasserstein ambiguity sets are widely used in the DRO literature due to their strong analytical and computational properties, including convex reformulations (e.g., [27, 28]). Unlike $\phi$-divergence balls, which may include statistically irrelevant distributions, Wasserstein balls account for distances between support elements and thus capture meaningful perturbations. This property prevents pathological distributions and leads to more reliable solutions [29].
The type-2 Wasserstein distance between two probability distributions $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{W})$ is defined as
@@
where $\mathcal{T}(\mathbb{P},\mathbb{Q})$ denotes the set of joint probability measures on $\mathcal{W}\times\mathcal{W}$ with marginals $\mathbb{P}$ and $\mathbb{Q}$. The optimization variable $\tau$ represents the transportation plan that redistributes probability mass from $\mathbb{P}$ to $\mathbb{Q}$. [^5] We use the Euclidean norm to measure the transportation cost $\|x-y\|$.
We define the ambiguity sets as Wasserstein balls centered at nominal distributions:
$\mathbb{D}_{j,t}\coloneqq\left\{\mathbb{P}_{j,t}\in\mathcal{P}(\real{n_{j}})\,|\,W_{2}(\mathbb{P}_{j,t},\hat{\mathbb{Q}}_{j,t})\leq\theta_{j}\right\}$
for $j\in\{w,v,x_{0}\}$, where $\theta_{j}$ denotes the Wasserstein ball radius, which specifies the level of robustness. [^6] Although we model the process and measurement noise uncertainties using separate ambiguity sets, the formulation can be extended to correlated noise by placing a single Wasserstein ball around their joint distribution. Extending the theory to such joint ambiguity sets is conceptually straightforward, but it leads to higher-dimensional optimization problems for filter design.
In general, computing $W_{2}$ explicitly is difficult. A useful surrogate is the Gelbrich distance. Specifically, the Gelbrich distance between $\mathbb{P}\in\mathcal{P}(\mathbb{R}^{n})$ and $\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{n})$ with mean vectors $\mu_{1},\mu_{2}\in\real{n}$ and covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}^{n}_{+}$ is defined as
@@
where
$\mathcal{B}(\Sigma_{1},\Sigma_{2})\coloneqq\sqrt{\operatorname{Tr}[\Sigma_{1}+\Sigma_{2}-2\big(\Sigma_{2}^{\frac{1}{2}}\Sigma_{1}\Sigma_{2}^{\frac{1}{2}}\big)^{\frac{1}{2}}]}$
is the Bures–Wasserstein distance.
The Gelbrich distance always provides a lower bound on the type-2 Wasserstein distance, i.e., $G(\mathbb{P},\mathbb{Q})\leq W_{2}(\mathbb{P},\mathbb{Q})$. Moreover, equality holds when both distributions are Gaussian, or more generally, elliptical with the same density generator [30].
@@ 3 Distributionally Robust Kalman Filters
We now develop tractable DR state estimators consistent with the noise-centric ambiguity model introduced in Section˜2. We first present a finite-horizon formulation whose recursion mirrors that of the KF under the least-favorable noise covariances. We then specialize to the time-invariant case and derive a steady-state DRKF that uses a single stationary SDP, resulting in a constant-gain filter with Kalman-level online complexity.
@@ 3.1 Finite-Horizon Case
We begin with assumptions on the nominal distributions that guarantee tractability of the DRSE problem and preserve the analytical structure of the classical KF, specifically the Gaussian and affine relationships that enable closed-form recursive updates.
@@
$(i)$ At the initial stage, the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
$\begin{split}\max_{\Sigma_{x,0}^{-},\Sigma_{v,0}}&\operatorname{Tr}[\Sigma_{x,0}^{-}-\Sigma_{x,0}^{-}C_{0}^{\top}(C_{0}\Sigma_{x,0}^{-}C_{0}^{\top}+\Sigma_{v,0})^{-1}C_{0}\Sigma_{x,0}^{-}]\\ \mbox{s.t.}\;&\mathrm{Tr}[\Sigma_{x,0}^{-}+\hat{\Sigma}_{x,0}^{-}-2((\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}}\Sigma_{x,0}^{-}(\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}})^{\frac{1}{2}}]\leq\theta_{x_{0}}^{2}\\ &\mathrm{Tr}[\Sigma_{v,0}+\hat{\Sigma}_{v,0}-2\big(\hat{\Sigma}_{v,0}^{\frac{1}{2}}\Sigma_{v,0}\hat{\Sigma}_{v,0}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}\\ &\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,0}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (4)
$(ii)$ For any $t>0$, fix the state distribution $\mathbb{P}_{x,t-1}=\mathcal{N}(\bar{x}_{t-1},\Sigma_{x,t-1})$. Then the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
@@
$\begin{split}\max_{\Sigma_{w,t-1},\Sigma_{v,t}}\;&\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2\big(\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\Sigma_{w,t-1}\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2\big(\hat{\Sigma}_{v,t}^{\frac{1}{2}}\Sigma_{v,t}\hat{\Sigma}_{v,t}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (5)
@@
$(iii)$ If $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ solve (4) and (5), respectively, then the maximum in (3) is attained by the Gaussian distributions $\mathbb{P}_{x,0}^{-,*}=\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-,*}),\mathbb{P}_{w,t}^{*}=\mathcal{N}(\hat{w}_{t},\Sigma_{w,t}^{*})$ and $\mathbb{P}_{v,t}^{*}=\mathcal{N}(\hat{v}_{t},\Sigma_{v,t}^{*})$ for each $t\geq 0$. Furthermore, at any time $t$, the optimal state estimator achieving the minimum in (3) is affine and given by
$\psi_{t}^{*}(\mathcal{Y}_{t})=\bar{x}^{-}_{t}+\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t}),$ (6)
with $\bar{x}^{-}_{t}=A_{t-1}\bar{x}_{t-1}+\hat{w}_{t-1},\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t}^{*}$, and $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*},\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$.
@@
The problems (4) and (5) define the stage-wise minimax interaction: the estimator selects $\psi_{t}$, while the adversary selects noise covariances within Wasserstein balls. Specifically, problem (4) yields the least-favorable prior and measurement covariances $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ at $t=0$, whereas (5) yields the least-favorable process and measurement covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ for $t\geq 1$. The corresponding optimal values represent the least-favorable posterior MSE at each stage.
@@
Lemma˜1 extends [17, Theorem 3.1] to the noise-centric setting by replacing ambiguity on the prior state (for $t>0$) with ambiguity on the process noise. Importantly, although the Wasserstein constraints are expressed in terms of first- and second-order moments, this does not restrict the adversary in a meaningful way under Assumption˜1. For Gaussian nominal distributions, the least-favorable distributions are Gaussian, and their worst-case behavior is fully characterized by their means and covariances. Furthermore, even without Gaussian nominal assumptions, (6) remains minimax optimal within the class of affine estimators [17].
@@
By adapting the structural argument used in [17, Lemma A.3], DRSE problem (3) admits at least one maximizer whose covariance matrices satisfy $\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}$, for $t=0$, and $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$ for all $t>0$ (see Appendix A). Because such a maximizer always exists, the explicit eigenvalue lower-bound constraints included in (4) and (5) are redundant: they do not alter the optimal value but simply restrict attention to maximizers that satisfy these bounds. In particular, the versions of (4) and (5) without these explicit bounds are equivalent to the original DRSE problem (3).
@@
In the inner maximization problems (4) and (5), the least-favorable distributions retain the nominal noise means. Any shift of a noise mean would introduce a systematic bias in the estimation error while also increasing the Wasserstein distance, making such a shift never advantageous for the adversary. Because of this trade-off, the ambiguity budget is used entirely to perturb the covariances, which is reflected in Lemma˜1 and Theorem˜1.
Both problems (4) and (5) are solvable for $\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$ since their objectives are continuous over compact feasible sets. Applying a standard Schur complement argument (e.g., [31, Appendix 5.5]), these problems reduce to SDPs that are efficiently solvable using off-the-shelf solvers.
@@
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,t}^{-},\Sigma_{x,t},\\ \Sigma_{w,t-1},\Sigma_{v,t},\\ Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,t}^{-}-\Sigma_{x,t}&\Sigma_{x,t}^{-}C_{t}^{\top}\\ C_{t}\Sigma_{x,t}^{-}&C_{t}\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w,t-1}&Y\\ Y^{\top}&\Sigma_{w,t-1}\end{bmatrix}\succeq 0,\;\begin{bmatrix}\hat{\Sigma}_{v,t}&Z\\ Z^{\top}&\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (7)
Similar results hold for $t=0$. Solving (7) offline for all $t$ yields the least-favorable covariances $(\Sigma_{w,t}^{*},\Sigma_{v,t}^{*})$. The next theorem shows that these covariances can be used directly in a Kalman-style recursion, resulting in a DR state estimator that is robust to distributional uncertainty.
@@
Under Assumption˜1, the minimax DR state estimate $\psi^{*}_{t}(\mathcal{Y}_{t})$ at any time $t$ coincides with the conditional mean $\bar{x}_{t}$ under the least-favorable distributions. Moreover, the least-favorable prior and posterior remain Gaussian: $\mathbb{P}_{x,t}^{-,*}=\mathcal{N}(\bar{x}^{-}_{t},\Sigma_{x,t}^{-})$, $\mathbb{P}_{x,t}^{*}=\mathcal{N}(\bar{x}_{t},\Sigma_{x,t})$. These distributions are recursively computed for $t=0,1,\dots$ as follows:
@@
    where $\Sigma_{v,t}^{*}$ is the maximizer of (5) for $t>0$ and (4) for $t=0$. Then, update $\bar{x}_{t}$ and $\Sigma_{x,t}$ based on the measurement $y_{t}$ as follows:
@@
    The initial values are $\bar{x}^{-}_{0}=\hat{x}_{0}^{-},\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$, where $\Sigma_{x,0}^{-,*}$ is the maximizer of (4) .
@@
    where $\Sigma_{w,t}^{*}$ is the maximizer of (5) for $t>0$.
Notably, the measurement update and state prediction equations (9)–(12) mirror those of the classical KF when the system is driven by the least-favorable distributions $\mathbb{P}_{x,0}^{-,*},\mathbb{P}_{w,t}^{*}$, and $\mathbb{P}_{v,t}^{*}$.

### v3

@@ 2.2 Wasserstein Ambiguity Sets
We adopt Wasserstein ambiguity sets due to their favorable analytical properties and their ability to exclude statistically irrelevant distributions (e.g., [21, 5]). The type-2 Wasserstein distance between two probability distributions $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{W})$ is defined as
@@
where $\mathcal{T}(\mathbb{P},\mathbb{Q})$ denotes the set of joint probability measures with marginals $\mathbb{P}$ and $\mathbb{Q}$. The optimization variable $\tau$ represents the transportation plan that redistributes probability mass from $\mathbb{P}$ to $\mathbb{Q}$. [^5] We use the Euclidean norm to measure the transportation cost $\|x-y\|$.
We define the ambiguity sets as Wasserstein balls centered at nominal distributions: for $j\in\{w,v,x_{0}\}$,
$\mathbb{D}_{j,t}\coloneqq\big\{\mathbb{P}_{j,t}\in\mathcal{P}(\real{n_{j}})\,|\,W_{2}(\mathbb{P}_{j,t},\hat{\mathbb{Q}}_{j,t})\leq\theta_{j}\big\},$
where $\theta_{j}$ specifies the robustness radius. [^6] The formulation can be extended to correlated noise by placing a single Wasserstein ball around the joint distribution, at the cost of higher-dimensional optimization problems.
In general, computing $W_{2}$ is difficult. A useful surrogate is the Gelbrich distance. Specifically, the Gelbrich distance between $\mathbb{P}\in\mathcal{P}(\real{n})$ and $\mathbb{Q}\in\mathcal{P}(\real{n})$ with mean vectors $\mu_{1},\mu_{2}\in\real{n}$ and covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}^{n}_{+}$ is defined as
@@
where $\mathcal{B}(\Sigma_{1},\Sigma_{2})\coloneqq\sqrt{\operatorname{Tr}[\Sigma_{1}+\Sigma_{2}-2\big(\Sigma_{2}^{\frac{1}{2}}\Sigma_{1}\Sigma_{2}^{\frac{1}{2}}\big)^{\frac{1}{2}}]}$ is the Bures–Wasserstein distance.
The Gelbrich distance always provides a lower bound on the type-2 Wasserstein distance, i.e., $G(\mathbb{P},\mathbb{Q})\leq W_{2}(\mathbb{P},\mathbb{Q})$. Moreover, equality holds when both distributions are Gaussian, or, more generally, elliptical with the same density generator [6].
@@ 3 Distributionally Robust Kalman Filters
We develop tractable DR state estimators consistent with the noise-centric ambiguity model in Section˜2. We first present a finite-horizon formulation whose recursion mirrors that of the classical KF when driven by least-favorable noise covariances. We then specialize to the time-invariant case and derive a steady-state DRKF obtained from a single stationary SDP, yielding a constant-gain filter with Kalman-level online complexity.
@@ 3.1 Finite-Horizon Case
We begin with a standard assumption on the nominal distributions that ensures tractability of the DRSE problem and preserves the affine–Gaussian structure underlying closed-form KF updates.
@@
$(i)$ At the initial stage, the following convex optimization problem has the same optimal value as (3), and any of its optimal solutions are optimal for (3):
@@
$\begin{split}\max_{\Sigma_{x,0}^{-},\Sigma_{v,0}}&\operatorname{Tr}[\Sigma_{x,0}^{-}-\Sigma_{x,0}^{-}C_{0}^{\top}(C_{0}\Sigma_{x,0}^{-}C_{0}^{\top}+\Sigma_{v,0})^{-1}C_{0}\Sigma_{x,0}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{x,0}^{-}+\hat{\Sigma}_{x,0}^{-}-2((\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}}\Sigma_{x,0}^{-}(\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}})^{\frac{1}{2}}]\leq\theta_{x_{0}}^{2}\\ &\operatorname{Tr}[\Sigma_{v,0}+\hat{\Sigma}_{v,0}-2\big(\hat{\Sigma}_{v,0}^{\frac{1}{2}}\Sigma_{v,0}\hat{\Sigma}_{v,0}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\;\Sigma_{v,0}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}\\ &\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,0}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (4)
$(ii)$ For any $t>0$, fix the state distribution $\mathbb{P}_{x,t-1}=\mathcal{N}(\bar{x}_{t-1},\Sigma_{x,t-1})$. Then the following convex optimization problem has the same optimal value as (3), and any of its optimal solutions are optimal for (3):
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{w,t-1},\\ \Sigma_{v,t},\Sigma_{x,t}^{-}\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2\big(\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\Sigma_{w,t-1}\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2\big(\hat{\Sigma}_{v,t}^{\frac{1}{2}}\Sigma_{v,t}\hat{\Sigma}_{v,t}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\;\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (5)
@@
$(iii)$ If $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ solve (4) and (5), respectively, then the maximum in (3) is attained by the Gaussian distributions $\mathbb{P}_{x,0}^{-,*}=\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-,*}),\mathbb{P}_{w,t-1}^{*}=\mathcal{N}(\hat{w}_{t-1},\Sigma_{w,t-1}^{*})$ and $\mathbb{P}_{v,t}^{*}=\mathcal{N}(\hat{v}_{t},\Sigma_{v,t}^{*})$ for each $t$. Furthermore, the minimax-optimal estimator is affine and given by
$\begin{split}\psi_{t}^{*}&(\mathcal{Y}_{t})=\bar{x}^{-}_{t}+\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t}),\end{split}$ (6)
with $\bar{x}^{-}_{t}=A_{t-1}\bar{x}_{t-1}+\hat{w}_{t-1},\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*},\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$.
@@
Problems (4) and (5) define the stage-wise minimax interaction: the estimator selects $\psi_{t}$, while the adversary selects noise covariances within Wasserstein balls. Problem (4) yields the least-favorable prior and measurement covariances $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ at $t=0$, whereas (5) yields the least-favorable process and measurement covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ for $t\geq 1$. The corresponding optimal values represent the least-favorable posterior MSE at each stage.
@@
Lemma˜1 extends [23, Thm. 3.1] to the noise-centric setting by replacing ambiguity on the prior state (for $t>0$) with ambiguity on the process noise. Although the Wasserstein constraints are equivalently expressible in terms of first and second moments, this does not meaningfully restrict the adversary under Assumption˜1. For Gaussian nominal distributions, the least-favorable distributions are Gaussian, and the worst-case behavior is fully characterized by their means and covariances. Furthermore, even without Gaussian nominal assumptions, (6) remains minimax optimal within the class of affine estimators [23].
@@
Although the ambiguity sets are defined as $W_{2}$ balls over full distributions, the maximization problems in (4) and (5) admit least-favorable distributions that preserve the nominal noise means. Indeed, for any fixed distribution, the minimizing estimator is the conditional mean, and the conditional MSE depends only on the posterior covariance through its trace. Mean shifts consume Wasserstein budget without increasing the worst-case MSE as effectively as perturbing the covariance. Hence, mean shifts are suboptimal for the adversary, and the ambiguity budget is used entirely to perturb covariances, as reflected in Lemma˜1 and Theorem˜1.
@@
By adapting the structural argument used in [23, Lemma A.3], the DRSE problem (3) admits at least one maximizer whose covariance matrices satisfy $\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}$, for $t=0$, and $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$ for all $t>0$ (see Appendix A). Because such a maximizer always exists, the explicit eigenvalue lower-bound constraints in (4) and (5) are redundant: they do not alter the optimal value but only restrict attention to maximizers that satisfy these bounds. Equivalently, removing these constraints yields problems with the same optimal value as (3).
Both problems (4) and (5) are solvable for $\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$ since their objectives are continuous over compact feasible sets. Applying the standard Schur complement argument, these problems reduce to SDPs that can be solved efficiently using off-the-shelf solvers.
@@
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1},\\ \Sigma_{v,t},Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,t}^{-}-\Sigma_{x,t}&\Sigma_{x,t}^{-}C_{t}^{\top}\\ C_{t}\Sigma_{x,t}^{-}&C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w,t-1}&Y\\ Y^{\top}&\Sigma_{w,t-1}\end{bmatrix}\succeq 0,\;\begin{bmatrix}\hat{\Sigma}_{v,t}&Z\\ Z^{\top}&\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\;\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (7)
Similar results hold for $t=0$. Solving (7) offline for all $t>0$ yields the least-favorable covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$. The next theorem shows that these covariances can be used directly in a Kalman-style recursion, resulting in a DR state estimator that is robust to distributional uncertainty.
@@
Under Assumption˜1, the minimax-optimal DR state estimate $\psi^{*}_{t}(\mathcal{Y}_{t})$ at any time $t$ coincides with the conditional mean $\bar{x}_{t}$ under the least-favorable distributions. Moreover, the least-favorable prior and posterior remain Gaussian: $\mathbb{P}_{x,t}^{-,*}=\mathcal{N}(\bar{x}^{-}_{t},\Sigma_{x,t}^{-})$, $\mathbb{P}_{x,t}^{*}=\mathcal{N}(\bar{x}_{t},\Sigma_{x,t})$. These distributions are recursively computed for $t=0,1,\dots$ as follows:
@@
    where $\Sigma_{v,t}^{*}$ is the maximizer of (5) for $t>0$ and (4) for $t=0$. Then, update $\bar{x}_{t}$ and $\Sigma_{x,t}$ as
@@
    The initial values are $\bar{x}^{-}_{0}=\hat{x}_{0}^{-},\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$, where $\Sigma_{x,0}^{-,*}$ is the maximizer of (4) at $t=0$.
@@
    where $\Sigma_{w,t}^{*}$ is the maximizer of (5) at time $t+1$.
Notably, the measurement update and state prediction equations (9)–(12) mirror those of the classical KF when the system is driven by the least-favorable distributions.